well if i have this kinda question what do i do
SAlly has twice as many Dimes as Nickels.the total value is &3.50.how many nickels does she has?

Found 2 solutions by checkley75, scott8148:
Answer by checkley75(3666)

.10*2X+.05*X=3.50
.2X+.05X=3.50
.25X=3.50
X=3.50/.25
X=14 NICKLES.
2*14=28 DIMES
PROOF
28*.10+14*.05=3.50
2.80+.70=3.50
3.50=3.50

Answer by scott8148(6628)

let x=nickels, so 2x=dimes ... multiply the number of coins by their value to find the total (working in cents is easiest)

5(x)+10(2x)=350 ... 25x=350 ... x=14 ... 2x=28